Instructor-created question Question Help A friend of yours bought a new sports car with a $4,000 down payment plus a $27,000 car loan that is financed at an interest rate of 0.50% per month for 60 months. After 2 years, the Blue Book value of her vehicle in the used-car marketplace is $16,000. a. Calculate the required monthly loan payment on the car. b. How much does your friend still owe on the car loan immediately after she makes her 24th payment? c. Compare your answer to Part (b) to $16,000. This situation is called being upside down. What can she do about it? d. If she decides to keep the car but pay $300 more each month, how many months wll it take her to payoff the remaining loan she owes? a. The required monthly payment is $ 522.1 (Round to the nearest cent.) b, Your friend still owes S 11777히 on the car loan. (Round to the nearest dolar.) c. What is the best thing to do in this situation? Choose the correct answer below *A. She should keep the car longer and hope the vehicle remains in good working order B. In case the mileage is high, the car is worth more than $16,000 and she should consider seling it )C. She should sell the car immediately D. She should stop paying out her loan as it is obviously unfair d. It will take her 15] months to payoff the remaining loan balance (Round-up to the nearest month) litk to select your answer(s) and then click Check Answer
Answer #1

a. $585

Explanation: =$27,000*0.5% =135 is interest per month on car loan

=$27,000 / 60 months = $450

The required monthly loan payment on the car is $585.

b. $21,060

Explanation: principal of every month calculated for 60 months is $450

So, including the interest $585*24 months = $14,040

The rest of car loan amount due is = $585*36 = $21,060

c. She should keep the car longer and hope the vehicle remains in good working order.

Explanation:

compared to part (b) answer to $16,000. The market place of the car is $16,000 but only the car of the principal amount due is $16,200 without interest.

If it calculated with interest of 0.5% for remaining 36 months is =135*36 =$4,860

So the overall due amount along with interest is $16,200 + $4,860

=$21,060.

If she sells the car as per the market value for $16,000 then the deal is not fair and her friend will not bear the excess of amount $5,060($585*36 = $21,060 – $16,000 actual pay her friend per market value)

So, the deal is not fair for her friend. She would better to drop the proposal of selling the car.

d. 22 months approximately

Explanation:

Suppose, If she decided to pay 300 more amount, it will clear the loan amount in 21.6 months means approximately 22 months.

Only principal amount per month is $450 + $300 excess amount she wants to pay more.

=$750

=$450*36 months

=$16,200

Let us calculate the increase of principal amount monthly

=$16,200/$750

=21.6 months or 22 months approximately.